OpenStudy (whtcutie):

I need help with this question.... The formula c=square root a^2+b^2 represents the length of the hypotenuse of a right triangle with side lengths a and b. solve the equation for b. show the work.

jhonyy9 (jhonyy9):

c = sqrt(a^2 +b^2) squared both sides and will get c^2 = a^2 +b^2 subtract from both sides a^2 c^2 -a^2 = b^2 make squareroot from both sides and will get b = sqrt(c^2 -a^2) hope helped

jhonyy9 (jhonyy9):

catie this is understandably ?

jhonyy9 (jhonyy9):

or do you have questions ?

OpenStudy (whtcutie):

no this is perfect thank you! but I do have another question that i need help on

OpenStudy (whtcutie):